On three consecutive passes a football team gains 5 yards, loses 17 yards and gains 33 yards. What number represents the total net yardage?

Answer by John10(245)

First, you gain 5 yards (+5)
Then------loses 17 yards (-17)
----------gain 33 yards (+ 33)
The net (sum) after three passes:
N = +5 + (-17) + (+33) = 5 - 17 + 33 = 21 yards
John10:)
P.S" If you like my help with any Algebra classes (Math 116 or 117 classes), please visit my site or email me. Thanks!